40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 23 10:16:44 2010, josephb, Update, CDS, Rerouted RFM around c1lsc, took RFM card out of c1lsc 
    Reply  Fri Jul 23 21:35:48 2010, rana, Update, CDS, Rerouted RFM around c1lsc, took RFM card out of c1lsc 
       Reply  Mon Jul 26 10:02:42 2010, josephb, Update, CDS, Rerouted RFM around c1lsc, took RFM card out of c1lsc 
Message ID: 3289     Entry time: Mon Jul 26 10:02:42 2010     In reply to: 3283
Author: josephb 
Type: Update 
Category: CDS 
Subject: Rerouted RFM around c1lsc, took RFM card out of c1lsc 

If you're refering to just the medm screen,  those can be restored from the SVN.  As we're moving to a new directory structure, starting with /opt/rtcds/caltech/c1/, the old LSC screens can all be put back in the /cvs/cds/caltech/medm/c1/lsc directory if desired.

The slow lsc aux crate, c1iscaux2, is still working, and those channels are still available.  I confirmed that one was still updating. As a quick test, I went to the SVN and pulled out the C1LSC_RFADJUST.adl file, renamed it to C1LSC_RFadjust.adl and placed it in /cvs/cds/caltech/medm/c1/lsc/, and checked it linked properly from the C1IOO_ModeCleaner.adl file.  I haven't touched the modulation depths, as I didn't want to mess with the mode cleaner, but if I get an OK, we can test that today and confirm that modulation depth control is still working.

Quote:

 I just realized that an unfortunate casualty of this LSC work was the deletion of the slow controls for the LSC which we still use (some sort of AUX processor). For example, the modulation

depth slider for the MC is now in an unknown state.

 

ELOG V3.1.3-